# The Key Physics of Ice Premelting

### v1

@@ XI How liquid is quasi-liquid?
However, it has been traditional in the study of the ice-vapor interface to refer to the surface disordered layer as the quasi-liquid layer. This is fine, but it puts the emphasis of the problem on how different this could be from the expected bulk liquid, suggesting enigmatic properties and an aura of mistery.
@@
The question then remains as to whether the small extent of liquid in between the surfaces is sufficient to exhibit a dynamic response similar to bulk water. Computer simulations are particularly insightful as they allow to monitor single molecule motion without perturbing the system. Unfortunately, measurements of self-diffusion coefficients are difficult to interpret, as they strongly depend on the location of molecules within the premelting film.Louden and Gezelter (2018); Kling, Kling, and Donadio (2018); Moreira et al. (2018) The results typically show that molecules close to the ice surface have a low diffusivity, while those at the liquid surface exhibit enhanced diffusion.Louden and Gezelter (2018) However, bulk properties cannot be expected for the three dimensional motion of the molecules, because of the strong confinement in the direction perpendicular to the film. Indeed, the motion in this direction can even become sub-diffusive or glassy-like, because of the locking and freezing of the molecules at the ice surface.Moreira et al. (2018) On the other hand, measuring the diffusion parallel to the interface for liquid-like molecules within the film, yields a rather different picture. In this case, the water molecules exhibit self-diffusion coefficients that are very close to those obtained for bulk undercooled water at the same temperature (Fig.15). This is sufficient for the film to display an almost bulk like rheological response when sheared. Indeed, simulations of the TIP4P/Ice model displayed a velocity pattern consistent with Couette flow, and a shear stress similar to that expected for bulk water, albeit with slip lengths in the scale of several nanometers (Fig.15).Baran et al. (2022) This results appear strongly in conflict with experimental measurements of viscosity coefficients estimated from the spreading of a liquid droplet siting on ice, which were found to be 20 to 200 times larger than for bulk water.Murata et al. (2015); Nagata et al. (2019); Sazaki et al. (2022) However, based on the description of wetting dynamics of premelting films, the motion of the contact line in a liquid drop is often dictated, not by the spontaneous flow of the liquid film, but by the growth rate of the ice surface underneath.Sibley et al. (2021) Therefore, the flow of a liquid film next to a droplet need not be a faithful proxy of the fluid properties. Indeed, the current understanding of nanofluidics is that there is no reason to expect such large viscosities in films of nanometer thickness.Bocquet and Charlaix (2010)
Refer to caption Refer to caption
@@
Figure 15: Rheology of premelting films at T=8T=-8 C as described by the TIP4P/Ice model. Results are shown for four different ice/substrate interfaces, with the substrate characterized by the contact angle of water droplets formed on the substrate. (a) Density profiles shown in either blue (for regions with liquid-like molecules as the majority phase) or black (for regions with solid-like molecules as the majority phase) continuous lines. The orange line shows the corresponding flow profiles resulting when the substrate slides at a speed of 5 m// on the ice surface. The dashed black lines show predictions for Couette flow, with slopes as given by the corresponding bulk water viscosity. (b) Diffusion coefficients calculated for the parallel motion of liquid-like molecules. The blue dashed line shows the result corresponding to bulk undercooled water. This is compared with parallel diffusion coefficients of water-like molecules in three different regions of the premelted layer as indicated by the dashed black lines of panel (a). Blue circles correspond to liquid-like molecules within the central region of the premelting film. Green squares correspond to molecules within the wall adsorption layer. The red diamonds correspond to the molecules in the ice adsorbed layer. Reproduced with permission from Baran et al.Baran et al. (2022) PNAS 119 e2209545119 (2022).
@@ XII Future Challenges
In between the sublimation line and the kinetic condensation line, growth proceeds in a quasi-stationary mode, whereby accretion of ice occurs while a stable premelting film remains in a steady state regime of constant thickness.Kuroda and Lacmann (1982); Kuroda (1983); Kuroda and Gonda (1990) This premelting film, even at sub-nanometer length-scales that occur at strong under-cooling has a profound impact on the crystal surface structure and promotes a sequence of structural surface phase transitions,Llombart, Noya, and MacDowell (2020) similar to those conjectured by Kuroda and Lacmann many years ago.Kuroda and Lacmann (1982) As a result, growth rates of the basal and prism faces crossover twice, and lead to the alternation of columnar and plate like hexagonal prisms described in the celebrated Nakaya diagram.Nakaya (1954); Furukawa and Wettlaufer (2007); Libbrecht (2022)
@@
Indeed, to a great extent the conclusions drawn in this perspective rely on computer simulations, which have been used both as a stand alone characterization technique, or as an essential tool for the interpretation of experiments.Sánchez et al. (2017); Rashmi and Paesani (2025); Berrens et al. (2025) However, simulations have some limitations. The most apparent for the study of ice premelting close to the triple point is the uncertainty in the determination of melting temperatures, which due to system size effects and unavoidable simulation artifacts (finite time step, truncation of interactions) is of the order of one Kelvin.Baran et al. (2024) On the other hand, at lower temperatures that are relevant for the study of incipient surface disorder, classical molecular dynamics could be unreliable at least in quantitative terms, as the strong libration frequencies of water molecules result in significant nuclear quantum effects.Paesani and Voth (2008) Be as it may, most of the studies of ice premelting by computer simulations have been performed with fixed point charge models,Abascal et al. (2005); Abascal and Vega (2005) which do not deal with environment dependent polarizability effects that could become very important at an interface. Another problem is that most fixed charged models predict a very small vapor pressure,Vega, Abascal, and Nezbeda (2006) which could be a matter of concern. Current advances in force field modeling are promising,Bore et al. (2022) and preliminary assessments of the extent of ice premelting by means of accurate first principle models such as MDPol seem to confirm the most significant features observed with classical force fields.Rashmi and Paesani (2025) Qualitative agreement from machine learned Density Functional Potentials would be also welcome, and some work in this direction has been published recently.Zeng and Gao (2024); Berrens et al. (2025) Another important issue is the study of ice premelting by computer simulations is the initial preparation of ice surfaces. Both basal and primary prism faces have a bilayer structure, and initial configurations can be prepared with a fully terminated or a half terminated bilayer. Some of the work reported in this study for the basal face was prepared with a half terminated bilayer,Llombart et al. (2019, 2020); Llombart, Noya, and MacDowell (2020) while that of the prism face was prepared with a fully terminated bilayer. Exactly what surface preparation is more relevant to the experimental situation is unclear, and further work in this direction is required to understand the actual self adsorption state of equilibrium premelting.
@@
Most of this perspective article has been concerned with ice premelting above 200 K. However, the ice surface at lower temperatures is also of relevance for the understanding of atmospheric and extraterrestrial physics, and poses interesting fundamental problems concerning the organization of dangling hydrogen bonds,Nojima et al. (2017), the competition between cubic and hexagonal ice,Hudait and Molinero (2016) and surface reconstruction.Fletcher (1992); Hong et al. (2024) Some time ago, Fletcher suggested that the ice surface at low temperature could be stabilized by adapting ordered hydrogen bond arrangements, in a so called Fletcher stripe phase.Fletcher (1992) This hypothesis has been explored in some studies for frozen realizations of bulk ice,Buch et al. (2008); Pan et al. (2010) but some work remains to understand how thermal agitation could change the picture, and what is the role of the conjectured order/disorder transition in the emergence of translational surface disorder.Sugimoto et al. (2019) A particularly interesting question is how the disordering of the Fletcher surface phase could be coupled to the underlying bulk hydrogen bond network, and whether the Fletcher phase is consistent with the hydrogen bond disordered network of ice Ih.Sugimoto et al. (2019); Berrens et al. (2025) Recent experimental work shows that the low temperature surface reconstruction of epitaxially grown patches could be actually far more complicated that previously envisaged.Hong et al. (2024)
@@
Another relevant conceptual problem refers to the role of electrostatic effects. In the study of aqueous interfaces, the interface potential always features an important contribution from electrostatic interactions. This have been discarded altogether in the current survey, under the assumption that water is neutral. However, in view of the shallow minimum of the interface potential, weak electrostatic effects could become significant at long range. e.g. is there a role of electrical effects in creating a second minimum of the interface potential…? Currently, both experimental,Dosch, Lied, and Bilgram (1996); Petrenko (1994) and theoretical reports,Watkins, VandeVondele, and Slater (2010); Slater and Michaelides (2019) suggest that the first few ice layers below the premelting film could exhibit significant Bjerrum defects, leading to an effective surface charge at the ice/water surface. This, together with charge accumulation from water’s self dissociation, might play a role in the properties of premelting films,Petrenko (1994) but an assessment with simple force fields is simply beyond reach. The formulation of a simple Poisson-Boltzmann theory is possible. But unfortunately, this requires imposing boundary conditions at the ice/water and water/vapor interfaces that appear currently not well understood. Indeed, the electrical nature of the water/vapor interface is a highly controversial issue, and even the definition and sign of the electric potential difference between the bulk phases appears to be unsettled.Becker, Loche, and Netz (2022) Yet, a far less discussed topic that is needed to interpret the role of electrostatics is the boundary condition for the ionic force at the solid phase: i.e. what is the pH of bulk ice? This was a problem that attracted the attention of Eigen,Eigen and Maeyer (1958) but does not seem to have been a matter of inquire in the literature ever since. Understanding both of these fundamental problems of chemical physics is interesting per se, and is absolutely necessary in order to assess whether electrostatic interactions promote or inhibit surface melting, and to what extent.
@@
Regarding the shallowness of the interface potential, it is interesting to recall the discrepancy of contact angle measurements of macroscopic and microscopic droplets.Huerre, Josserand, and Séon (2025) If the contact angle of 12 measured for large droplets were to be confirmed as the correct macroscopic equilibrium contact angle, the depth of the free energy minimum of the interface potential would increase significantly, and some of the problems mentioned above could be remedied. This begs the question: Could the discrepancy be related to line tension effects? Close to a first order wetting transition, line tensions are expected to be large and positive, and promote the spreading of small droplets.Indekeu (1994) A large line tension also implies that the condensation of liquid droplets is an activated process requiring a degree of supersaturation for the droplet to form,Navascues and Tarazona (1981); Widom (1995) in line with the findings of the Hokaido team.Murata et al. (2016) However, it is clear that the early stages of droplet deposition on the ice surface do not correspond to full thermodynamic equilibrium, so one cannot rule out that the discrepancy is related to pinning of the contact line by frozen water in some extremely complex out of equilibrium process.
@@
Beyond surface ice premelting, it has been amply verified that liquid-like layers can also be significant at interfaces,Dash (1999); Engemann et al. (2004); Liljeblad, Furó, and Tyrode (2017) and some recent computer simulations confirm this for a wide range of substrates.Rönneberg et al. (2020); Baran et al. (2022); Cui et al. (2024); Atila, Sukhomlinov, and Müser (2025) However, there is still a large room for understanding the role of a substrate on ice premelting, and what interfacial characteristics enhance or inhibit growth of quasi-liquid layers.Baran, Llombart, and MacDowell (2025)
@@ Acknowledgements.
I would like to express my most sincere gratitude to my collaborators and friends who have helped me understand some of the intricate physics of ice premelting. First of all I am particularly indebted to former PhD students and PostDocs Jorge Benet, Pablo Llombart, Juan Luengo, Fernando Izquierdo-Ruiz and Lukasz Baran, who have performed all of the computational work, and stimulated me with many illuminating discussions, and particularly, to Lukasz Baran, for invaluable discussions, help and ongoing collaboration. I am also greatly indebted to my colleagues, Eva G. Noya, David Sibley, Andrew Archer, Mathias Böstrom and Eduardo Sanz for their willingness to embark in our collaboration and for many helpful discussions. I also wish to express my gratitude for discussions and support to Valeria Molinero, Keneth Libbrecht, Kenichiro Murata, Dan Gezelter, Davide Donnadio, Gilles Demange, Cristophe Josserand, Menno Demmenie, Emily Aesenath-Smith and many others. The research discussed here was partially funded by the Agencia Estatal de Investigación (Ministerio de Ciencia y Economía) under grant PID2023-151751NB-I00.

### v2

@@ XI How liquid is quasi-liquid?
However, it has been traditional in the study of the ice-vapor interface to refer to the surface disordered layer as the quasi-liquid layer. This is fine, but it puts the emphasis of the problem on how different this could be from the expected bulk liquid, suggesting enigmatic properties and an aura of mystery
@@
The question then remains as to whether the small extent of liquid in between the surfaces is sufficient to exhibit a dynamic response similar to bulk water. Computer simulations are particularly insightful as they allow to monitor single molecule motion without perturbing the system. Unfortunately, measurements of self-diffusion coefficients are difficult to interpret, as they strongly depend on the location of molecules within the premelting film.Louden and Gezelter (2018); Kling, Kling, and Donadio (2018); Moreira et al. (2018) The results typically show that molecules close to the ice/water surface have a low diffusivity, while those at the water/vapor surface exhibit enhanced diffusion.Louden and Gezelter (2018) However, bulk properties cannot be expected for the three dimensional motion of the molecules, because of the strong confinement in the direction perpendicular to the film. Indeed, the motion in this direction can even become sub-diffusive or glassy-like, because of the locking and freezing of the molecules at the ice/water surface.Moreira et al. (2018) On the other hand, measuring the diffusion parallel to the interface for liquid-like molecules within the film, yields a rather different picture. In this case, the water molecules exhibit self-diffusion coefficients that are very close to those obtained for bulk undercooled water at the same temperature (Fig.15). This is sufficient for the film to display an almost bulk like rheological response when sheared. Indeed, simulations of the TIP4P/Ice model displayed a velocity pattern consistent with Couette flow, and a shear stress similar to that expected for bulk water, albeit with slip lengths in the scale of several nanometers (Fig.15).Baran et al. (2022) This results appear strongly in conflict with experimental measurements of viscosity coefficients estimated from the spreading of a liquid droplet siting on ice, which were found to be 20 to 200 times larger than for bulk water.Murata et al. (2015); Nagata et al. (2019); Sazaki et al. (2022) However, based on the description of wetting dynamics of premelting films, the motion of the contact line in a liquid drop is often dictated, not by the spontaneous flow of the liquid film, but by the growth rate of the ice surface underneath.Sibley et al. (2021) Therefore, the flow of a liquid film next to a droplet need not be a faithful proxy of the fluid properties. Indeed, the current understanding of nanofluidics is that there is no reason to expect such large viscosities in films of nanometer thickness.Bocquet and Charlaix (2010)
[…]
@@
Figure 15: Rheology of premelting films at T=8T=-8 C as described by the TIP4P/Ice model. Results are shown for ice at the interface with four substrates with differing water contact angle . (a) Density profiles shown in either blue (for regions with liquid-like molecules as the majority phase) or black (for regions with solid-like molecules as the majority phase) continuous lines. The orange line shows the corresponding flow profiles resulting when the substrate slides at a speed of 5 m/s/s on the ice surface. The dashed black lines show predictions for Couette flow, with slopes as given by the corresponding bulk water viscosity. (b) Diffusion coefficients calculated for the parallel motion of liquid-like molecules. The blue dashed line shows the result corresponding to bulk undercooled water. This is compared with parallel diffusion coefficients of water-like molecules in three different regions of the premelted layer as indicated by the dashed black lines of panel (a). Blue circles correspond to liquid-like molecules within the central region of the premelting film. Green squares correspond to molecules within the wall adsorption layer. The red diamonds correspond to the molecules in the ice adsorbed layer. Reproduced with permission from Baran et al.Baran et al. (2022) PNAS 119 e2209545119 (2022).
@@ XII Future Challenges
In between the sublimation line and the kinetic condensation line, growth proceeds in a quasi-stationary mode, whereby accretion of ice occurs while a stable premelting film remains in a steady state regime of constant thickness.Kuroda and Lacmann (1982); Kuroda (1983); Kuroda and Gonda (1990) This premelting film, even at sub-nanometer length-scales that occur at strong under-cooling, has a profound impact on the crystal surface structure and promotes a sequence of structural surface phase transitions,Llombart, Noya, and MacDowell (2020) similar to those conjectured by Kuroda and Lacmann many years ago.Kuroda and Lacmann (1982) As a result, growth rates of the basal and prism faces crossover twice, and lead to the alternation of columnar and plate like hexagonal prisms described in the celebrated Nakaya diagram.Nakaya (1954); Furukawa and Wettlaufer (2007); Libbrecht (2022)
@@
Indeed, to a great extent the conclusions drawn in this perspective rely on computer simulations, which have been used both as a stand alone characterization technique, or as an essential tool for the interpretation of experiments.Sánchez et al. (2017); Rashmi and Paesani (2025); Berrens et al. (2025) However, simulations have some limitations. The most apparent for the study of ice premelting close to the triple point is the uncertainty in the determination of melting temperatures, which due to system size effects and unavoidable simulation artifacts (finite time step, truncation of interactions) is of the order of one Kelvin.Baran et al. (2024) On the other hand, at lower temperatures that are relevant for the study of incipient surface disorder, classical molecular dynamics could be unreliable at least in quantitative terms, as the strong libration frequencies of water molecules result in significant nuclear quantum effects.Paesani and Voth (2008) Be as it may, most of the studies of ice premelting by computer simulations have been performed with fixed point charge models,Abascal et al. (2005); Abascal and Vega (2005) which do not deal with environment dependent polarizability effects that could become very important at an interface. Another problem is that most fixed charged models predict a very small vapor pressure,Vega, Abascal, and Nezbeda (2006) which could be a matter of concern. Current advances in force field modeling are promising,Bore et al. (2022) and preliminary assessments of the extent of ice premelting by means of accurate first principle models such as MDPol seem to confirm the most significant features observed with classical force fields.Rashmi and Paesani (2025) Qualitative agreement from machine learned Density Functional Potentials would be also welcome, and some work in this direction has been published recently.Zeng and Gao (2024); Berrens et al. (2025) Another important issue in the study of ice premelting with computer simulations is the initial preparation of ice surfaces. Both basal and primary prism faces have a bilayer structure, and initial configurations can be prepared with a fully terminated or a half terminated bilayer. Some of the work reported in this study for the basal face was prepared with a half terminated bilayer,Llombart et al. (2019, 2020); Llombart, Noya, and MacDowell (2020) while that of the prism face was prepared with a fully terminated bilayer. Exactly what surface preparation is more relevant to the experimental situation is unclear, and further work in this direction is required to understand the actual self adsorption state of equilibrium premelting.
@@
Most of this perspective article has been concerned with ice premelting above 200 K. However, the ice surface at lower temperatures is also of relevance for the understanding of atmospheric and extraterrestrial physics, and poses interesting fundamental problems concerning the organization of dangling hydrogen bonds,Nojima et al. (2017), the competition between cubic and hexagonal ice,Hudait and Molinero (2016) and surface reconstruction.Fletcher (1992); Hong et al. (2024) Some time ago, Fletcher suggested that the ice surface at low temperature could be stabilized by adapting ordered hydrogen bond arrangements, in a so called Fletcher stripe phase.Fletcher (1992) This hypothesis has been explored in some studies for frozen realizations of bulk ice,Buch et al. (2008); Pan et al. (2010) but some work remains to understand how thermal agitation could change the picture, and what is the role of the conjectured order/disorder transition in the emergence of translational surface disorder.Sugimoto et al. (2019) A particularly interesting question is how the disordering of the Fletcher surface phase could be coupled to the underlying bulk hydrogen bond network, and whether the Fletcher phase is consistent with the hydrogen bond disordered network of ice Ih.Sugimoto et al. (2019); Berrens et al. (2025) Recent experimental work shows that the low temperature surface reconstruction of epitaxially grown patches could be actually far more complicated than previously envisaged.Hong et al. (2024)
@@
Another relevant conceptual problem refers to the role of electrostatic effects. In the study of aqueous interfaces, the interface potential always features an important contribution from electrostatic interactions. This have been discarded altogether in the current survey, under the assumption that water is neutral. However, in view of the shallow minimum of the interface potential, weak electrostatic effects could become significant at long range. e.g. is there a role of electrical effects in creating a second minimum of the interface potential…? Currently, both experimental,Dosch, Lied, and Bilgram (1996); Petrenko (1994) and theoretical reports,Watkins, VandeVondele, and Slater (2010); Slater and Michaelides (2019) suggest that the first few ice layers below the premelting film could exhibit significant Bjerrum defects, leading to an effective surface charge at the ice/water surface. This, together with charge accumulation from water’s self dissociation, might play a role in the properties of premelting films,Petrenko (1994) but an assessment with simple force fields is simply beyond reach. The formulation of a simple Poisson-Boltzmann theory is possible. But unfortunately, this requires imposing boundary conditions at the ice/water and water/vapor interfaces that appear currently not well understood. Indeed, the electrical nature of the water/vapor interface is a highly controversial issue, and even the definition and sign of the electric potential difference between the bulk phases appears to be unsettled.Becker, Loche, and Netz (2022) Yet, a far less discussed topic that is needed to interpret the role of electrostatics is the boundary condition for the ionic force at the solid phase: i.e. what is the pH of bulk ice? This was a problem that attracted the attention of Eigen,Eigen and Maeyer (1958) but does not seem to have been a matter of inquiry in the literature ever since. Understanding both of these fundamental problems of chemical physics is interesting per se, and is absolutely necessary in order to assess whether electrostatic interactions promote or inhibit surface melting, and to what extent.
@@
Regarding the shallowness of the interface potential, it is interesting to recall the discrepancy of contact angle measurements of macroscopic and microscopic droplets.Huerre, Josserand, and Séon (2025) If the contact angle of 12 measured for large droplets were to be confirmed as the correct macroscopic equilibrium contact angle, the depth of the free energy minimum of the interface potential would increase significantly, and some of the problems mentioned above could be remedied. This begs the question: Could the discrepancy be related to line tension effects? Unfortunately, assessing line tensions close to a wetting transition is challenging, and even the sign of the line tension remains a controversial issue.Indekeu (1994) However, notice that assuming a negative line tension would imply that the contact angle decreases as the droplets become smaller,Navascues and Tarazona (1981); Widom (1995) bringing into agreement the discrepancies between microscopic and macroscopic realizations of water droplets on ice. Yet, it is clear that the early stages of droplet deposition on the ice surface do not correspond to full thermodynamic equilibrium, so one cannot rule out that the discrepancy is related to pinning of the contact line by frozen water in some extremely complex out of equilibrium process.
@@
Beyond surface ice premelting, it has been amply verified that liquid-like layers can also be significant at interfaces,Dash (1999); Engemann et al. (2004); Liljeblad, Furó, and Tyrode (2017) and some recent computer simulations confirm this for a wide range of substrates.Rönneberg et al. (2020); Baran et al. (2022); Cui et al. (2024); Atila, Sukhomlinov, and Müser (2025) However, there is still a large room for understanding the role of a substrate on ice premelting, and what interfacial characteristics enhance or inhibit growth of quasi-liquid layers.Baran, Llombart, and MacDowell (2025) These interfacially premelted layers apparently exhibit close to bulk like properties,Baran et al. (2022) and can provide for a lubrication effect in ice sliding.Baran et al. (2022); Baran and MacDowell (2024)
@@ Acknowledgements.
I would like to express my most sincere gratitude to my collaborators and friends who have helped me understand some of the intricate physics of ice premelting. First of all I am particularly indebted to former PhD students and PostDocs Jorge Benet, Pablo Llombart, Ramón Bergua, Juan Luengo, Fernando Izquierdo-Ruiz and Lukasz Baran, who have performed all of the computational work, and stimulated me with many illuminating discussions, and particularly, to Lukasz Baran, for invaluable discussions, help and ongoing collaboration. I am also greatly indebted to my colleagues, Eva G. Noya, David Sibley, Andrew Archer, Mathias Boström and Eduardo Sanz for their willingness to embark in our collaboration and for many helpful discussions. I also wish to express my gratitude for discussions and support to Valeria Molinero, Keneth Libbrecht, Kenichiro Murata, Dan Gezelter, Davide Donnadio, Gilles Demange, Cristophe Josserand, Menno Demmenie, Emily Aesenath-Smith and many others. The research discussed here was partially funded by the Agencia Estatal de Investigación (Ministerio de Ciencia y Economía) and Fondo Europeo de Desarrollo Regional (FEDER) under grant PID2023-151751NB-I00.